OpenStudy (anonymous):

The random variable X is said to be stochastically larger than the random variable Y if P(X>z)>or= P(Y>z) for all real z with strict inequality holding for at least one z value. Show that this requires that the cdfs enjoy the following property: Fx(z)

OpenStudy (zarkon):

\[F_{X}(z)=P(X\le x),F_{Y}(z)=P(Y\le x)\] if \[P(X>z)\ge P(Y>z)\] then \[-P(X>z)\le -P(Y>z)\] \[1-P(X>z)\le 1-P(Y>z)\] \[P(X\le z)\le P(Y>\le z)\] \[F_{X}(z)\le F_{Y}(z)\]